Fermi National Laboratory


Accelerator Update for September 23, 2005 - September 26, 2005

Friday September 23
The day shift began with Operations monitoring store 4399 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with the MiniBooNE and NuMI experiments taking beam.

Operations began preparing the Recycler to stash antiprotons from Pbar at 9:59 AM. The Recycler stash ended at 11:05 AM and Pbar resumed stacking.

Pbar experts began a reverse proton study at 1:03 PM.

Main Injector (MI) experts began a damper study at 1:19 PM.

The Pbar study ended around 5:15 PM.

A Linac expert began working on fluctuating intensities at 6:17 PM.

Saturday September 24
The midnight shift began with Operations monitoring store 4399 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

Operations terminated store 4399 at 12:08 AM.

Operations put the Tevatron (TeV) into shot setup at 12:29 AM.

Operations established combination store 4400 at 2:52 AM with an initial luminosity of 124.69E30. Pbar resumed stacking.

The Recycler began stashing antiprotons at 5:48 AM.

At 6:33 AM, Operations reported that the last transfer of antiprotons from Pbar to the Recycler failed to inject. The MI beam permit was down due to a MECAR trip. Operators rebooted MECAR. Pbar resumed stacking.

Operations established beam to MCenter around 7:11 AM.

At 9:13 AM, a MI RF station (MIRF14) tripped off. It reset okay.

The Recycler began stashing at 10:42 AM. The stash ended and Pbar resumed stacking at 11:53 AM. MCenter resumed taking beam.

At 12:10 PM, Operations reported that they were having difficulty resuming normal stacking. An operator called in a Pbar expert.

At 12:18 PM, MIRF14 tripped off for the first of four times in two and a half hours. It reset each time.

A Pbar expert arrived around 1 PM and began investigating the stacking problem. He had it resolved and Pbar back to normal stacking in about forty-five minutes.

The Recycler began preparing to stash at 5:43 PM. The stash ended and Pbar and MCenter resumed normal operations at 6:46 PM.

At 10:16 PM, MIRF14 tripped off. It reset okay.

At 11:30 PM, Operations reported that a TeV RF station (TRF2) was off for 39 seconds and then turned itself back on. There was little effect on luminosity.

Sunday September 25
The midnight shift began with Operations monitoring store 4400 and the stash, and with MiniBooNE and NuMI taking beam.

At 5:37 AM, Operations reported having more stacking problems. Operators found a frequency problem and made adjustments.

Operations terminated store 4400 around 8 AM.

Operations put the TeV into shot setup at 8:47 AM.

Operations established combination store 4401 at 11:20 AM with an initial luminosity of 131.02E30. This store set a NEW LUMINOSITY RECORD. Pbar and MCenter resumed normal operations.

The Recycler began stashing at 1:57 PM. The stash ended at 2:54 PM. Pbar and MCenter resumed normal operations.

Recycler stash: 7:05 PM - 8:11 PM. Pbar and MCenter resumed normal operations.

Monday September 26
The midnight shift began with Operations monitoring store 4401 and the stash, and with MiniBooNE, NuMI, and MCenter taking beam.

Recycler stash: 12:14 AM - 12:50 AM. Pbar and MCenter resumed normal operations.

The Plans for Monday
The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar experts to conduct a reverse protons study during the day shift.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/26/2005   email Fermilab